ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 42A05, 42A16 Taghavi, Mohsen, Shiraz University, Iran Exact values of the Autocorrelation coefficients of the Rudin-Shapiro polynomials The maximal order of the Autocorrelation coefficients of a Rudin-Shapiro polynomial of “length” L is dominated by CL^{0.73} and the lower bound for those non-zero coefficients is CL^{0.5}, where C is an absolute constant. Using the square representation of their Fourier coefficients, we introduce some new techniques to find the exact values of all such coefficients.

1. M. Taghavi, An estimate on the Correlation coefficients of the Rudin-Shapiro polynomials, Iranian Journal of Science and Technology, Vol. 20, No. 2, 1996.

2. M. Taghavi, Upper bounds for the Autocorrelation coefficients of the Rudin-Shapiro Polynomials, Korean Journal of Computational and Applied Mathematics, Vol. 4, No. 1, 1997.

3. M. Taghavi, An extremal problem applied to the Rudin-Shapiro polynomials, Korean Journal of Computational and Applied Mathematics, Vol. 5, No. 1, 1998.